how many liters of a 90% acid solution must be mixed with a 15% acid solution to get 600 liters of a 80% solution?
Answer by Alan3354(69443)

how many liters of a 90% acid solution must be mixed with a 15% acid solution to get 600 liters of a 80% solution?
=======================
n = liters of 90%
f = liters of 15%
----
n + f = 600 ---------- total solution
90n + 15f = 600*80---- total acid
---
etc